OpenStudy (anonymous):

Ed has 100 more dollars than robert. after ed spends 20 dollars on groceries, he now has 5 times as much money as robert. how much money does robert have. the answer the book has is 20

OpenStudy (saifoo.khan):

Let x=amount that Robert has Then Ed has x+100 So, our equation to solve is: (x+100)-20=5x get rid of parens x+100-20=5x x+80=5x subtract x from each side x-x+80=5x-x collect like terms 80=4x divide each side by 4 x=20 dollars-------------------amount that Robert has x+100=20+100=120 dollars ----amount that Ed has CK 120-20=5*20 100=100